# Can LLMs Replace Manual Annotation of Software Engineering Artifacts?

### v1

@@ I Introduction
Given that developer effectiveness heavily depends on suitable tools and processes, researchers are seeking to innovate by improving and automating parts of software engineering. To mention just a few example, there are now many research efforts [1, 2] aimed at code summarization, i.e., techniques that generate a natural language summary for a given piece of code, at detecting bugs and other issues in a program [3], and at determining whether a warning produced by a static analysis tool is actually worth addressing [4].
@@
The usefulness of such tools and techniques ultimately is a question of human cognition. For example, a developer might read an English summary of a method, e.g., in a docstring, and then decide whether and how to use that method. As another example, a developer might look at a warning produced by a static analysis tool, and then decide whether to fix it or not. Because measuring human usefulness is difficult, many research efforts rely on proxy metrics. For the example of code summarization, a summary might rate a reasonable BLEU score, indicating at least some level of similarity to the original, human-written summary. However, even if the BLEU score is good, how can we know if such a summary is clear, precise, relevant to the input code, and human-comprehensible?
@@
To answer such questions, researchers strongly advocate for, and adopt human subject-based evaluations, e.g., when evaluating code summarization techniques [5, 6]. Naturally, this evaluation approach is not just for tools that perform code summarization. Such evaluations are becoming increasingly common in many different settings; indeed in most settings where some aspect of software production is automated using a supporting tool, whether software artifacts are being generated or analyzed, the output from the supporting tool will be used by a human. Since the content & form of the output may affect the ability of the human to properly use the output, it may be sometimes necessary to evaluate the tool using a human-subject study.
@@
Human-subject studies in software engineering are costly. For external validity, such studies require both a representative sample of developers and a representative sample of relevant artifact samples. Each sample may need to be evaluated by multiple humans, to get more stable results. For example, Haque et al. [5], hired professional developers, at \$60 per hour, to rate a total of 420 human and machine-generated code summaries, with three ratings for each sample to gauge inter-rater agreements. Due to the high costs, such evaluations are sometimes done using free or low-cost participants, such as students, which obviously carries some threats, as to whether the results would generalize to professional developers. Yet, even performing a human-subject study with students can be quite time-consuming. In other words, the very reason that motivates researching new tools and techniques in software engineering – the high cost of developers – is challenging the proper evaluation of such tools and techniques.
@@
Given the advances in large language models (LLMs), and their ability to match human performance in a range of tasks [7, 8, 9], the question naturally arises: Can we use LLMs to reduce the cost of human-subject studies in software engineering? Providing a positive answer, even if only for a subset of the effort involved in such studies, could have a huge impact on the practice of evaluations in software engineering. Moreover, improving our understanding of when and how human study participants could be safely replaced by LLMs is an important first step toward mixed human-LLM evaluations.
@@
The above question has arisen in other disciplines, e.g., psychology, linguistics, and medicine [10, 11, 12, 13]. There are several reasons why the software engineering domain is particularly interesting for this exploration, at this time. First, the relative cost: It’s quite costly to recruit representative subjects, whereas LLMs are cheap enough to run hundreds of queries for the cost of using a single human subject for an hour. Second, LLMs are well-trained on software artifacts, both code and code-related natural language, have proven to be useful to a wide range of software engineering tasks, and are now in routine use in the software industry [14, 15, 7, 16]. Third, advances in prompt engineering, such as the impressive abilities of LLMs at in-context learning [17, 18], suggest that with a few illustrative examples, LLMs can execute fairly complex software engineering tasks. Finally, software artifacts are arguably more complex than natural language artifacts: they can require deep knowledge of both application domain and programming; also, they comprise multiple formal & informal elements (e.g, code, summary, warnings).
@@
We study the possibility of substituting human subjects with LLMs when annotating software engineering artifacts. We study ten tasks from five datasets created by prior work, covering research ranging from requirements engineering to reasoning about static analysis warnings. We apply six SOTA LLMs, including both open- and closed-source models, to these tasks, and then compare the LLMs’ responses to those from human subjects. Our findings include:
@@
  • On some tasks, LLMs agree with human subjects about as much as human subjects agree with each other. For example, when judging whether an identifier name and the value it refers to are consistent [19], we observe a mean human-to-human inter-rater agreement of 0.52, and a mean human-to-model agreement of 0.49. This results suggests that LLMs can be a meaningful replacement for humans on some annotation tasks.
@@
  • To help researchers decide whether an LLM may be suitable for a specific annotation task, we investigate the correlation between model-to-model agreement and the ability of LLMs to meaningfully replace humans. We find a strong positive correlation (Spearman correlation of 0.65), which suggests that model-to-model agreement can serve as a relatively cheap-to-obtain proxy to decide whether to involve LLMs in an evaluation.
@@
  • Some examples are easier to annotate than others. For a given example, can we tell whether an LLM can substitute for a human? We find that the LLM’s confidence (its output probability) helps select examples that do not necessarily require human annotators. For example, for the task of judging whether a code snippet and its summary are similar, delegating 50% of the effort performed by one annotator to an LLM (selected based on LLM confidence), does not lead to a statistically significant change of the overall inter-rater agreement.
@@
  • We are the first to study whether LLMs can replace manual annotation on software engineering artifacts.
@@
  • We propose a methods for selecting which tasks are suitable for an LLM-augmented evaluation and which samples can be safely delegated to an LLM.
@@ II Background
In software engineering research, evaluations of tools and techniques often rely on qualitative annotations of software artifacts (e.g., code, documentation, bug reports, or newly generated research artifacts [20]). These annotations provide insight into the effectiveness of innovations and their potential impact on developer productivity or software quality. Traditionally, many such evaluations involve human subjects, which requires multiple annotators to ensure consistency and reliability in the data [21].
Inter-rater reliability is vital because it provides a benchmark for the objectivity of the annotations [22]. It is commonly measured using statistical metrics such as Cohen’s $\kappa$ [23] or Krippendorff’s $\alpha$ [24], which measure agreement among annotators by taking into account the possibility of agreement occurring by chance. High values indicate a strong agreement among annotators, suggesting that the annotations are reliable and that the findings are robust.
When annotations are based on a single annotator, there is a significant risk that the findings may reflect personal biases and subjective interpretations, which could lead to inaccuracies and limit the generalizability of the results [25]. In contrast, multiple annotators working independently can achieve a shared understanding, which ensures that the annotations accurately capture the underlying phenomena being studied. However, achieving high inter-rater reliability is effort-intensive and time-consuming, often involving substantial manual work. This challenge is compounded by the variability in human interpretation, which can lead to inconsistencies in the data.
LLMs hold promise to adress these challenges by potentially providing more consistent annotations. Properly trained LLMs could reduce the biases introduced by human annotators and enable the scaling of qualitative analysis to much larger datasets. The integration of LLMs into the annotation process could open up new research opportunities by making it feasible to analyze datasets that were previously too large or complex to annotate manually. In addition, software engineering data is dynamic and constantly evolving, requiring tools that can quickly adapt to new information. LLMs, with their ability to learn from and adapt to new data, could provide up-to-date annotations, complementing the work of human annotators.
To effectively incorporate LLMs into qualitative annotation, it is essential to establish a structured process. This includes deciding when and how to use LLMs, integrating them into existing workflows, distributing tasks between humans and LLMs, and determining the degree of manual work required. This paper addresses these considerations through four research questions, focusing on annotations with pre-existing categories derived from related work or developed in an initial research phase. Future work will explore the potential of LLMs for annotating qualitative data without pre-existing categories (i.e., open coding) within a software engineering context.
@@ III Methodology
In this section, we describe the tasks and datasets, the models under consideration, our research questions, and the methodology to answer these questions. We tried to use the exact same instructions for the models that were used in the original study.
@@ III-A Tasks & Datasets
We select five datasets from previous work, which in total contain ten annotation tasks. The datasets and tasks are selected to represent a diverse range of software engineering research.
Code Summarization Automatic code summarization is an active area aiming to generate helpful summaries of code [3]. One survey-based study reports that 80% of respondents found code comment generation tools useful [26]. 78% agree that these tools help them understand the source code, especially helping code readability in projects with few comments. Several metrics are used to evaluate code summary quality. All metrics aim to indirectly measure human perception of quality, but have limitations. But what criteria are important to evaluate human perception? Haque et al. [5] considered four criteria: accuracy, adequacy, conciseness, and similarity—that should be considered when evaluating code summary quality.
Haque et al. [5] recruited experienced programmers via Upwork, paying them USD60/hr (the applicable market rate). The programmers were presented with 210 functions (from LeClair et al [27]) along with associated human-generated, and model-generated summaries (420 in total). Subjects responded to four questions, related to each of the criteria above. Each participant rated a subset of the 420 samples [^1] https://github.com/similarityMetrics/similarityMetrics, and the authors ensured that each sample was seen by at least three raters.
To evaluate an LLM on this task, we prompt the model with the same guidelines and questions as in the original study. That is, the LLM receives the functions and associated comments, and is tasked to rate its agreement with four statements:
@@
Name-Value Inconsistencies Variable names are crucial for code understanding [19]. If a variable name does not match the value stored in it, we have an undesirable “name-value inconsistency”.
Patra and Pradel [19] user study had 11 participants, and created a dataset evaluating name-value consistency [^3] https://github.com/sola-st/Nalin. This dataset includes 40 samples rated by all 11 raters. The raters assigned a Likert score from 1 to 5, where 1 indicates difficulty in understanding (or the presence of name-value inconsistency), and 5 indicates ease of understanding. Suppose there is a variable named “name” to which a float, 2.5, is assigned. This is difficult for developers to understand because the expected value was a string. All annotators rated it as difficult to understand. On the contrary, another variable “done”, assigned the boolean value False, is easy to understand and was rated as easy to understand by most annotators. We challenge the LLM to perform the same tasks, again using few-shot prompting.
Causality Causal knowledge supports reasoning about requirements dependencies, particularly for tasks such as test-case construction. Causal relations (e.g., ‘If event 1, then event 2’) are common in system behavior. However, extracting causality from natural language documents is difficult. Fischbach et al. [29] introduced a dataset, built using 6 raters to develop and evaluate a tool for extracting causal dependencies. [^4] https://github.com/fischJan/CiRA The dataset is quite large, with over 10,000 samples. We randomly select 1,000 samples where at least two raters assessed each sentence. The ratings are binary: 1 indicates the presence of a causal relation, and 0 indicates its absence. We then ask the LLM to perform the same annotation task as the original participants based on a few-shot prompt.
@@
Static Analysis Warnings Automatic static analysis tools, (e.g. FindBugs), can find coding errors, but do often raise false alarms. Pruning false alarms and presenting only (or mostly) actionable warnings to developers can be beneficial. A dataset by Kang et al[4] provides human annotations on the task of determining whether a particular code change effectively addresses a static analysis warning. The dataset has 1,306 samples that cover different warning categories. [^6] https://github.com/soarsmu/SA_retrospective/tree/main Two annotators assigned each sample one of three labels: open, closed, or unknown. A static analysis warning is considered “closed” if it was reported in a previous software revision but not in a subsequent reference revision, indicating that the problematic code was altered or removed. Conversely, a warning is labeled as “open” if it appears in both the current testing revision and a later reference revision, suggesting that the warning was not actionable or it was ignored by the developers. Lastly, a warning is marked as “unknown” if the file containing the warning was deleted or modified in a (possibly unrelated) way in the reference revision, making it difficult to confirm whether the warning was actionable. Given the extensive metadata, the prompts for this problem are the longest. Due to rate limitations imposed by some models (e.g., Claude and Gemini) and the need to manage costs, we use 200 samples, chosen uniformly at random, for our experiments. For each of these samples, we ask LLMs to perform the annotation task.
@@ III-C Research Questions
Our goal is to study whether LLM prompting can subsitute for human annotation labour, for artifact annotation tasks in SE. Specifically, we want to understand how annotator (inter-rater) agreement changes when substituting human annotations with model annotations. In the first research question, we examine all three categories of inter-rater agreement: human-human, human-model, and model-model. Specifically, if human-model inter-rater agreement is similar to that of human-human agreement, it may be possible to substitute (some) human ratings with model ratings. We will also examine the inter-model agreement and how it changes across datasets.
@@
In Section IV-A we discuss the findings of human-human, human-model and model-model inter-rater agreement. Depending on the dataset and annotationt task, we find varying levels of agreement. This raises the question of how one might decide whether a particular task is amenable to replacing human effort with an LLM.
@@ III-D Evaluation Methodology
For RQ1: With each task, we record a) number of annotators and b) samples rated by each annotator. Note that an annotator may not have annotated all the samples in the dataset but rather a subset. [^7] For example, Haque et al used six annotators, but each sample only received three annotations. After identifying the samples annotated by each human, we compute the inter-rater agreement metric Krippendorff’s $\alpha$ for each human-human, human-model, and model-model pair. To clarify, if a human (P1) has rated 5 samples {s1, s2, s3, s4, and s5}, and another human (P2) has rated 5 samples {s3, s4, s5, s6, s7}, we consider only {s3, s4, s5} for computing Krippendorff’s $\alpha$ for P1 and P2. There are other inter-rater agreement metrics, but Krippendorff’s $\alpha$ is applicable to any number of annotators, each assigning one value to one unit of analysis, to incomplete (missing) data, to any number of values available for annotating a variable, and to different kinds of annotations (binary, nominal, ordinal, interval, ratio, etc.). Given the diversity of the datasets we study, this metric allows us to uniformly apply inter-rater agreement to all datasets.

### v2

@@ Abstract
§§footnotetext: Majority of the work was done when the author was a postdoctoral scholar at UC Davis.
@@ I Introduction
Given that developer effectiveness heavily depends on good tools and processes, researchers constantly seek more and better automation in these areas. To mention just a few examples, there are now many research efforts [1, 2] aimed, e.g., at automated code summarization, (i.e., techniques that generate a natural language summary for a given piece of code), at detecting bugs and other issues in a program [3], and at determining whether a warning produced by a static analysis tool is actually worth addressing [4].
@@
But the value of such innovations ultimately is dependent on human judgment and practice. For example, a developer might read a generated English summary of a method, and then decide whether and how to use that method. As another example, a developer might look at a warning produced by a static analysis tool, and then decide whether to fix it or not. Because measuring human usefulness is difficult, many research efforts rely on proxy metrics. For the example of code summarization, a summary might rate a reasonable BLEU score, indicating at least some level of similarity to the original, human-written summary. However, even given a summary with a high BLEU score, can we be sure that it is clear, precise, relevant to the input code, and human-comprehensible?
@@
To answer such questions, researchers often use human subject-based evaluations, e.g., when evaluating code summarization techniques [5, 6]. Besides code summarization, such evaluations are becoming increasingly common in many different settings; indeed in most cases where a tool is used to automate some aspect of software development, the output from the tool will be used by a human. Since the content and form of the output may affect how well a developer can use it, a human-subject evaluation is often vital.
@@
But human-subject studies in software engineering are costly. For external validity, such studies demand both a representative sample of developers and a representative sample of relevant artifact samples. Each sample may need to be evaluated by multiple humans, to get more stable results. For example, Haque et al. [5], hired professional developers, at \$60 per hour, to rate a total of 420 human and machine-generated code summaries, with three ratings for each sample to gauge inter-rater agreements. Due to the high costs, such evaluations are sometimes done using free or low-cost participants, such as students, which carries risks of the results not generalizing to professional developers. Yet, even performing a human-subject study with students can be quite time-consuming. Thus, the main motivation for researching new tools and techniques in software engineering, viz. the high cost of developers, is also a hurdle to the proper evaluation of such tools and techniques.
@@
Given the ability of advanced large language models (LLMs), to rival human performance in a range of tasks [7, 8, 9], the question naturally arises: Can we use LLMs to reduce the cost of human-subject studies in software engineering? If LLMs can help, even if only partially, this could impact the practice of evaluation studies in software engineering. Thus, we, ask: When, and how, can human subject responses be safely replaced by LLMs, in a mixed human-LLM evaluation scenario?
@@
This question has arisen in other disciplines, e.g., psychology, linguistics, and medicine [10, 11, 12, 13]. There are several reasons why the software engineering domain is particularly interesting for this exploration: First, relative cost: LLMs can run hundreds of queries for the cost of using a single human subject for an hour. Second, LLMs now perform well on a wide range of software engineering tasks, and are routinely used in industry [14, 15, 7, 16]. Third, the impressive capabilities of LLMs at in-context learning [17, 18], suggest that with a few illustrative examples, LLMs can execute fairly complex software engineering tasks. Finally, software artifacts are arguably more complex than natural language artifacts: they can require deep knowledge of both application domain and programming; also, they comprise multiple formal & informal elements (e.g, code, summary, warnings).
@@
This paper studies the possibility of substituting human subjects with LLMs when annotating software engineering artifacts. We study ten tasks from five datasets created by prior work, covering research ranging from requirements engineering to reasoning about static analysis warnings. We apply six state-of-the-art LLMs, including both open- and closed-source models, to these tasks, and then compare the LLMs’ responses to those from human subjects.
We focus on tasks whose subjective, nuanced and context-dependent nature have traditionally required human judgment. This subjectivity is evident in the fact prior studies reveal considerable inter-rater disagreement. Figure 1 shows two examples from code summarization, where human subjects are asked to rate the accuracy of generated summaries. In Figure 1(a), all the three raters and the GPT-4 model rated “Strongly agree”. However, other examples are inherently hard for humans, such as Figure 1(b), where the three raters give different ratings: “Agree”, “Strongly Disagree”, and “Disagree”, and GPT-4 agrees with the first rater. Our study asks: Can powerful LLMs (partially) substitute for humans on inherently subjective annotation tasks?. Our findings are:
(a) Sample where human raters agreed
(b) A sample where humans disagreed
Figure 1: Annotation tasks with different difficulty level.
@@
  • On some tasks, LLMs agree with human subjects about as much as human subjects agree with each other. For example, when judging whether an identifier name and the value it refers to are consistent [19], we observe a mean human-human inter-rater agreement of 0.52, and a mean human-model agreement of 0.49. This results suggests that LLMs can sometimes be a meaningful replacement for humans on some annotation tasks. But when, exactly?
@@
  • To help researchers decide whether an LLM may be suitable for a specific annotation task, we investigate the correlation between model-model agreement, which can be measured programmatically, and the ability of LLMs to meaningfully replace humans. We find model-model agreement strongly correlates with human-model agreement, suggesting that model-model agreement can be used to decide whether to involve LLMs in an evaluation.
@@
  • Some examples are easier to annotate than others. For a given example, can an LLM substitute for a human? We find that the LLM’s confidence (its output probability) helps select examples that do not necessarily require human annotators. For example, for the task of judging whether a code snippet and its summary are similar, delegating 50% of the effort performed by one annotator to an LLM (selected based on LLM confidence), does not lead to a statistically significant change of the overall inter-rater agreement.
@@
  • We offer a first study as to whether LLMs can replace human annotation on software engineering artifacts.
@@
  • We propose a method for selecting which tasks are suitable for an LLM-augmented evaluation and which samples can be safely delegated to an LLM.
@@ II Background
In software engineering research, human-subject annotations provide insight into the potential impact of innovations on productivity or quality [20]; but the nuanced, subjective nature of these evaluations often requires multiple human raters to ensure consistency and reliability in the data [21].
Inter-rater reliability provides a vital gauge for the objectivity of the annotations [22]. It is commonly measured using statistical metrics such as Cohen’s $\kappa$ [23] or Krippendorff’s $\alpha$ [24], which measure agreement among annotators by considering the possibility of pure-chance. Higher values suggest a strong, reliable, and robust agreement among annotators.
Annotations based on a single annotator carry a risk of personal biases and subjective interpretations, and lead to inaccurate, non-generalizable results [25]. In contrast, multiple annotators working independently can reveal an inherently shared understanding, which indicates that the annotations accurately capture the underlying phenomena being studied. However, achieving high inter-rater reliability is effortful, time-consuming, and challenging; human developer time is costly, and they often disagree!
LLMs, since they are trained on very large corpora, could help these challenges by potentially providing more consistent annotations which normatively reflect “most” developers. They could reduce the biases introduced by human annotators, LLM-based automation of annotations could make it feasible to analyze datasets that were previously too large or complex to annotate manually. In addition, software engineering data is dynamic and constantly evolving, requiring tools that can quickly adapt to new information. LLMs, with their ability to learn from and adapt to new data, could provide up-to-date annotations, complementing the work of human annotators.
To effectively incorporate LLMs into qualitative annotation, it is essential to establish a structured process. This includes deciding when and how to use LLMs, integrating them into existing workflows, distributing tasks between humans and LLMs, and determining the degree of manual work required. This paper addresses these considerations through four research questions (See subsection III-C), focusing on annotations with pre-existing categories derived from related work or developed in an initial research phase. Future work will explore the potential of LLMs for annotating qualitative data without pre-existing categories (i.e., open coding) within a software engineering context.
@@ III Methodology
We now describe the tasks, datasets, the models under consideration, our research questions, and the methodology to answer these questions.
@@ III-A Tasks & Datasets
We select five datasets from previous work, which together include 10 human-annotation tasks. For some datasets, multiple tasks were assigned to the raters. For example, in the code summarization dataset, raters were given four tasks: rating accuracy, adequacy, similarity, and conciseness. Similarly, for the semantic similarity dataset, raters were assigned three different rating tasks. This is why we have five datasets or major tasks but a total of 10 human-annotation tasks. In general, we ask LLMs to perform annotations previously done by humans; we give the LLMs the same instructions as given to humans. The datasets and tasks are selected to represent a diverse range of software engineering research.
Automatic Code Summarization This is an active area, aiming to generate helpful summaries of code [3]. One survey-based study reports that 80% of respondents found code comment generation tools useful [26]. 78% agree that these tools help them understand the source code, especially helping code readability in projects with few comments. Several metrics are used to evaluate code summary quality. All metrics aim to indirectly measure human perception of quality, but have limitations. But what criteria are important to evaluate human perception? Haque et al. [5] considered four criteria: accuracy, adequacy, conciseness, and similarity—that should be considered when evaluating code summary quality.
Haque et al. [5] recruited experienced programmers via Upwork, paying them USD60/hr (the applicable market rate). The programmers were presented with 210 functions (from LeClair et al [27]) along with associated human-generated, and model-generated summaries (420 in total). Subjects responded to four questions, related to each of the criteria above. Each participant rated a subset of the 420 samples [^1] https://github.com/similarityMetrics/similarityMetrics, while ensuring that each sample was rated by at least three people.
To elicit LLM responses, we prompt the model with the same guidelines and questions as in the original study. That is, the LLM receives the functions and associated comments, and is tasked to rate its agreement with four statements:
@@
Name-Value Inconsistencies Variable names are crucial for code understanding [19]. If a variable name does not match the value stored in it, we have an undesirable “name-value inconsistency”. Patra and Pradel [19] user study had 11 participants, and created a dataset evaluating name-value consistency [^3] https://github.com/sola-st/Nalin. This dataset includes 40 samples rated by all 11 raters. The raters assigned a Likert score from 1 to 5, where 1 indicates difficulty in understanding (or the presence of name-value inconsistency), and 5 indicates ease of understanding. Suppose there is a variable named name to which a float, 2.5, is assigned. This is confusing for developers, since the expected value is a string. All annotators rated it as difficult to understand. On the contrary, another variable done, assigned the boolean value False, is easy to understand and was rated as easy to understand by most annotators. We challenge the LLM to perform the same tasks, again with a few-shot prompt.
Causality Causal knowledge supports reasoning about requirements dependencies, particularly for tasks such as test-case construction. Causal relations (e.g., ‘If event 1, then event 2’) are common in system behavior. However, extracting causality from natural language documents is difficult. Fischbach et al. [29] introduced a dataset, built using 6 raters to develop and evaluate a tool for extracting causal dependencies. [^4] https://github.com/fischJan/CiRA The dataset is quite large, with over 10,000 samples. We randomly select 1,000 samples where at least two raters assessed each sentence. The ratings are binary: 1 indicates the presence of a causal relation, and 0 indicates its absence. We elicit such annotations from LLM using a few-shot prompt.
@@
Static Analysis Warnings Automatic static analysis tools, (e.g. FindBugs), can find coding errors, but can raise false alarms. Pruning false alarms and presenting only (or mostly) actionable warnings to developers can be beneficial. A dataset by Kang et al[4] provides human annotations on the task of determining whether a particular code change effectively addresses a static analysis warning. The dataset has 1,306 samples that cover different warning categories. [^6] https://github.com/soarsmu/SA_retrospective/tree/main Two annotators assigned each sample one of three labels: open, closed, or unknown. A static analysis warning is considered “closed” if it was reported in a previous software revision but not in a subsequent reference revision, indicating that the problematic code was altered or removed. Conversely, a warning is labeled as “open” if it appears in both the current testing revision and a later reference revision, suggesting that the warning was not actionable or it was ignored by the developers. Lastly, a warning is marked as “unknown” if the file containing the warning was deleted or modified in a (possibly unrelated) way in the reference revision, making it difficult to confirm whether the warning was actionable. We use diff to present the changes made to the repository. Given the extensive metadata, the prompts for this problem are the longest. Due to rate limitations imposed by some models (e.g., Claude and Gemini) and the need to manage costs, we use 200 samples, chosen uniformly at random, for our experiments. For each of these samples, we ask LLMs to perform the annotation task.
@@ III-C Research Questions
We seek to study whether LLM prompting can substitute for human annotation labour, for artifact annotation tasks in SE. Specifically, we study how annotator (inter-rater) agreement changes when replacing human annotations with automated LLM annotations. In the first research question, we examine all three categories of inter-rater agreement: human-human, human-model, and model-model. Specifically, if human-model inter-rater agreement is similar to that of human-human agreement, it may be possible to substitute (some) human ratings with model ratings. We will also examine the inter-model agreement and how it changes across datasets.
@@
In Section IV-A we discuss the findings of human-human, human-model and model-model inter-rater agreement. Depending on the dataset and annotation task, we find varying levels of agreement. This raises the question of how one might decide whether a particular task is amenable to replacing human effort with an LLM.
@@ III-D Evaluation Methodology
For RQ1: With each task, we record a) number of annotators and b) samples rated by each annotator. Note that an annotator may not have annotated all the samples in the dataset but rather a subset. [^7] For example, Haque et al used six annotators, but each sample only received three annotations. After identifying the samples annotated by each human, we compute the inter-rater agreement metric Krippendorff’s $\alpha$ for each human-human, human-model, and model-model pair. To clarify, if a human (P1) has rated 5 samples {s1, s2, s3, s4, and s5}, and another human (P2) has rated 5 samples {s3, s4, s5, s6, s7}, we consider only {s3, s4, s5} for computing Krippendorff’s $\alpha$ for P1 and P2. Other inter-rater agreement metrics exist; but Krippendorff’s $\alpha$ is applicable to any number of annotators, each assigning one value to one unit of analysis, to incomplete (missing) data, to any number of values available for annotating a variable, and to different kinds of annotations (binary, nominal, ordinal, interval, ratio, etc.). Given the diversity of the datasets we study, this metric allows us to uniformly apply inter-rater agreement to all datasets.